OpenStudy (bekkah323):

Simplify, leaving as little as possible inside a solute-value signs. I x^2/-y I and I-8x^2/ 2x I

OpenStudy (bekkah323):

sorry i meant absolute- value

OpenStudy (zzr0ck3r):

x^2 >= allways so |x^2| is redundant so take it out x^2/|-y|

OpenStudy (bekkah323):

what do you mean x^2 = always?

OpenStudy (zzr0ck3r):

same logic on the second (4x^2)||x|

OpenStudy (zzr0ck3r):

x^2 is always greater than or equal to zero

OpenStudy (zzr0ck3r):

that should say (4x^2)/|x|

OpenStudy (bekkah323):

the answer in my book for the second one was 4IxI? can you do more to that problem? and also how did you go from 8 to 4?

OpenStudy (bekkah323):

nvm i get the 4 to 8

OpenStudy (zzr0ck3r):

I-8x^2/ 2x I = |-8/2|*|x^2/x| = |-4||x| = 4|x|